# Language Models Learn Rare Phenomena from Less Rare Phenomena: The Case of the Missing AANNs

### v2

@@ 1.1 Motivation and Prior Work
There are different levels of memorization, though, requiring different levels of abstraction. Consider the aann construction: “a beautiful five days in Texas” (Solt, 2007; Keenan, 2013; Dalrymple and King, 2019), which is rarer than the default “five beautiful days in Texas”. A model that strictly memorizes this phrase might come to know that “a beautiful five days in Texas” is grammatical but have no idea that “a beautiful four days in Texas” is grammatical if it never appeared in its training. A model that generalizes just a bit more might know that “a beautiful five days in New York” is also grammatical by generalizing that any U.S. state can fill the slot. Knowing that “an astonishing 200 pages” is acceptable requires generalization beyond mere lexical items. And knowing that “a blue five pencils” is not acceptable (because colors are “stubbornly distributive”, Schwarzschild 2011) requires yet more knowledge. Even for an idealized learner, it is difficult to precisely formulate how these kinds of generlizations emerge.
@@ 1.2 Summary of findings
First, we find BabyLM-trained LMs to successfully generalize to novel instances of the aann construction. Performance unsurprisingly drops for LMs that were trained without being exposed to even a single aann during training, but perhaps surprisingly, not by all that much—they are substantially above chance. This suggests that certain items present in the training data might give rise to LMs’ non-trivial performance in judging acceptability of aanns. This finding is further strengthened by the fact that LMs trained on counterfactual variants of the aann—e.g., anan and naan, obtained by shuffling word order and are far less likely to share structure with training data items—are unable to generalize to those constructions as well as they do to aanns (one which they have not seen at all).
@@ 2.5 Scoring and Accuracy
and $\blacktriangle$ under the aann training condition are cases where training data was polluted by randomly replacing 300 aanns with anans and naans, respectively, in order to assess the impact of an imperfect aann detection system. The dashed line represents chance performance (6.25%) and the dot-dashed line represents accuracies for 2- and 4-gram LMs trained on BabyLM. Accuracies for GPT-2-XL (Radford et al., 2019) and Llama-2-7B (Touvron et al., 2023) are computed using log-probabilities, since unigram frequencies were unavailable for these LMs’ corpora.
Importantly, we train the unigram estimator for a given corpus using the same tokenizer used to train our autoregressive LMs on that corpus. We use slor in lieu of the usual normalized log-probability measure, ensuring that the comparison between two models cannot be explained simply by the difference in unigram frequencies due to our manipulations. Log-probabilities were computed using minicons (Misra, 2022). An instance within our test set is considered to be correct iff the slor values of the well-formed construction is greater than that for all four corrupted instances. The accuracy, then, is the proportion of correct instances within the test set. Since this involves four pairwise comparisons, chance performance is 6.25%.
@@ 3 Experiment 1: LMs learn about aanns without having seen a single instance
For comparison to larger, state-of-the-art LMs, we test Llama-2-7B (Touvron et al., 2023) and GPT-2 XL (Radford et al., 2019) on the aanns. They got 83% and 78%, respectively. As a comparison to shallower LMs, we tested on 2-and 4-gram LMs trained on BabyLM and found both got 0% accuracy, strongly suggesting that the observed results are not due to n-gram statistics.
@@
Given that BabyLM-trained LMs learn the aann construction, how well would an LM generalize to aanns without having seen a single positive instance? To this end, we compare accuracies in the previous experiment to that obtained by LMs trained on BabyLM with our 2,448 detected aanns removed (i.e., no aann). From Fig. 2, we find LMs trained with the no aann condition to achieve an average accuracy of 47%, which is a noticeable drop compared to the 70% obtained by the LMs trained on the complete BabyLM corpus, but importantly 40.75 points above chance performance (and, as we show below, well above comparable baselines with other constructions). This is a non-trivial result, since it suggests that LMs can learn the acceptability of a construction without having seen a single positive occurrence, which indicates that there exist systematic patterns in the corpus driving this generalization behavior.
@@
While we focused on overall structural properties of aanns, there are also idiosyncrasies to the construction that arise from lexical semantic constraints. For instance, in many aann sentences, people prefer quantitative adjectives such as mere and hefty to qualitative ones such as beautiful (Mahowald, 2023; Solt, 2007) and find “stubbornly distributive” adjectives (“*a blue five pencils”) completely unacceptable (Schwarzschild, 2011). Insofar as our models learn aanns, we also should expect them to learn these lexical constraints. To test this, we compared LMs’ slors to human acceptability judgments on all 3.4k instances in Mahowald’s data across different adjective and noun classes. We found LMs trained on the original, unmodified BabyLM corpus to pattern similarly to humans in their preference for lexical constraints affecting aanns. Interestingly, these patterns were unchanged for LMs trained with the no aann condition, conforming to our predictions. For instance, as seen in Fig. 3, both our models share the human preference for quantitative and qualitative adjectives in the aann, compared to stubbornly distributive adjectives. More detailed results on lexical constraints can be found in App. E and we hypothesize that our broader set of results extends to include learning of lexical constraints on the construction.
@@ 5 Experiment 3: The Role of Variability
We hypothesize that instances of aanns that provide natural evidence of greater open-slot variability—i.e. evidence that many different adjectives, numerals, and nouns can go in their respective positions in the aann construction—would lead LMs to assign greater likelihood to unseen aanns. On the other hand, LMs that encounter only a restricted set of instances might be more conservative in extending the coverage of possible aanns to novel combinations of the slot-fillers. To test this, we divided our set of 2,448 aann-containing utterances in the BabyLM corpus into two roughly equal subsets—one that contained aanns which were restricted in which tokens occur in a particular slot (low variability), and the other where the aanns showed more variability in those slots. We obtain these subsets by performing a median split based on the number of unique occurrences in a target slot(s), which resulted in a set of 1224 low and high variability instances. We repeated this for all three open slots (adjective/numeral/noun) jointly as well as those slots individually—i.e., 4 different types of target slots and 2 conditions each (low vs. high variability). Details about the slot fillers and examples from each set are provided in App. F. We then trained LMs on the BabyLM corpus containing utterances involving either of these two cases. Fig. 5 shows the resulting average slors obtained from this experiment, along with those obtained by LMs trained on unablated BabyLM and the no aann conditions.

### v3

@@ 1.1 Motivation and Prior Work
There are different levels of memorization, though, requiring different levels of abstraction. Consider the aann construction: “a beautiful five days in Texas” (Solt, 2007; Keenan, 2013; Dalrymple and King, 2019), which is rarer than the default “five beautiful days in Texas”. A model that strictly memorizes this phrase might come to know that “a beautiful five days in Texas” is grammatical but has no idea that “a beautiful four days in Texas” is grammatical if it never appeared in its training. A model that generalizes just a bit more might know that “a beautiful five days in New York” is also grammatical by generalizing that any U.S. state can fill the slot. Knowing that “an astonishing 200 pages” is acceptable requires generalization beyond mere lexical items. And knowing that “a blue five pencils” is not acceptable (because colors are “stubbornly distributive”, Schwarzschild 2011) requires yet more knowledge. Even for an idealized learner, it is difficult to precisely formulate how these kinds of generalizations emerge.
@@ 1.2 Summary of findings
First, we find BabyLM-trained LMs to successfully generalize to novel instances of the aann construction. Performance unsurprisingly drops for LMs that were trained without being exposed to even a single aann during training, but perhaps surprisingly, not by all that much—they are well above chance. This suggests that certain items present in the training data might give rise to LMs’ non-trivial performance in judging acceptability of aanns. This finding is further strengthened by the fact that LMs trained on counterfactual variants of the aann—e.g., anan and naan, obtained by shuffling word order and are far less likely to share structure with training data items—are unable to generalize to those constructions as well as they do to aanns (one which they have not seen at all).
@@ 2.5 Scoring and Accuracy
[…]
Importantly, we train the unigram estimator for a given corpus using the same tokenizer used to train our autoregressive LMs on that corpus. We use slor in lieu of the usual normalized log-probability measure, ensuring that the comparison between two models cannot be explained simply by the difference in unigram frequencies due to our manipulations. Log-probabilities were computed using minicons (Misra, 2022). An instance within our test set is considered to be correct iff the slor value of the well-formed construction is greater than that for all four corrupted instances. The accuracy, then, is the proportion of correct instances within the test set. Since this involves checking how often the LM prefers the target surface form out of five options, chance performance is 20%.
@@ 3 Experiment 1: LMs learn about aanns without having seen a single instance
For comparison to larger, state-of-the-art LMs, we test Llama-2-7B (Touvron et al., 2023) and GPT-2 XL (Radford et al., 2019) on the aanns. They got 83% and 78%, respectively. As a comparison to shallower LMs, we tested on 4-gram LMs trained on BabyLM and found them to get much lower accuracies (41%), suggesting that the observed results are beyond n-gram statistics.
@@
Given that BabyLM-trained LMs learn the aann construction, how well would an LM generalize to aanns without having seen a single positive instance? To this end, we compare accuracies in the previous experiment to that obtained by LMs trained on BabyLM with our 2,448 detected aanns removed (i.e., no aann). From Fig. 2, we find LMs trained with the no aann condition to achieve an average accuracy of 47%, which is a noticeable drop compared to the 70% obtained by the LMs trained on the complete BabyLM corpus, but importantly 27 points above chance performance (and, as we show below, well above comparable baselines with other constructions). This is a non-trivial result, since it suggests that LMs can learn the acceptability of a construction without having seen a single positive occurrence, which indicates that there exist systematic patterns in the corpus driving this generalization behavior.
@@
While we focused on overall structural properties of aanns, there are also idiosyncrasies to the construction that arise from lexical semantic constraints. For instance, people prefer quantitative adjectives such as mere and hefty to qualitative ones such as beautiful (Mahowald, 2023; Solt, 2007) and find “stubbornly distributive” adjectives (“*a blue five pencils”) completely unacceptable (Schwarzschild, 2011). Insofar as our models learn aanns, we also should expect them to learn these lexical constraints. To test this, we compared LMs’ slors to human acceptability judgments on all 3.4k instances in Mahowald’s data across different adjective and noun classes. We found LMs trained on the original, unmodified BabyLM corpus to pattern similarly to humans in their preference for lexical constraints affecting aanns. Interestingly, these patterns were unchanged for LMs trained with the no aann condition, conforming to our predictions. For instance, as seen in Fig. 3, both our models share the human preference for quantitative and qualitative adjectives in the aann, compared to stubbornly distributive adjectives. More detailed results on lexical constraints can be found in App. E and we hypothesize that our broader set of results extends to include learning of lexical constraints on the construction.
@@ 5 Experiment 3: The Role of Variability
We hypothesize that instances of aanns that provide natural evidence of greater open-slot variability—i.e. evidence that many different adjectives, numerals, and nouns can go in their respective positions in the aann construction—would lead LMs to assign greater likelihood to unseen aanns. On the other hand, LMs that encounter only a restricted set of instances might be more conservative in extending the coverage of possible aanns to novel combinations of the slot-fillers. To test this, we divided our set of 2,448 aann-containing utterances in the BabyLM corpus into two roughly equal subsets—one that contained aanns which were restricted in which tokens occur in a particular slot (low variability), and the other where the aanns showed more variability in those slots. We obtain these subsets by performing a median split based on the number of unique occurrences in a target slot(s), which resulted in a set of 1224 low and high variability instances. We repeated this for all three open slots (adjective/numeral/noun) jointly as well as those slots individually—i.e., 4 different types of target slots and 2 conditions each (low vs. high variability). Details about the slot fillers and examples from each set are provided in App. F. We then trained LMs on the BabyLM corpus containing utterances involving either of these two cases. Fig. 5 shows the average slors obtained from this experiment, along with those obtained by LMs trained on unablated BabyLM and the no aann conditions.
@@ 8 Acknowledgments
9 Corrigendum
The original version of this paper had two bugs affecting baselines, which have been fixed in the present version. 1) Because of a bug in how we handled the output from KenLM, the 4-gram slor calculation used log base 10 for the numerator and base $e$ for the denominator which affected accuracies for 4-gram models. This makes the 4-gram baselines stronger in Experiment 1 than they appeared to be, although the LM still outperforms the 4-gram baseline in the critical condition in which AANNs are removed. In this version, we have updated the 4-gram baseline results to fix the bug. We believe the 4-gram baseline should be seen as stronger than initially represented. But we do not believe this significantly alters the overall conclusions of the paper. We also updated Experiment 2 to fix the 4-gram baselines; those results are qualitatively unchanged. We thank Forrest Davis for bringing this bug to our attention. 2) Chance performance, as shown in Fig. 2, was incorrectly calculated as 6.25% ($\frac{1}{2}^{4}$) due to an erroneous independence assumption—it should be 20% as in this version. This does not qualitatively affect any conclusions.